increase or 7-9. Including activity coefficients, find [Hg3] in saturated Hg,Br2 in 0.001 00 M KBr.
Answer #1

First calculate ionic strength of KBr solution.

We have, phpGtkzZ2.png= 1/2 Sum ( C i Z i2 )

Where C i is concentration of i th species and Z i is its charge.

Charge on K is +1 and on Br is -1.

[K+] = [Br - ] = 0.00100 M

\therefore The ionic strength is phpKoWcaf.png = 1/2 [ [K+] (+1) 2 + [Br - ] (-1) 2 ]

phpKoWcaf.png = 1/2 [ 0.00100 *1 + 0.00100 *1 ]

phpKoWcaf.png = 0.00100 M

For phpQO2xkK.png= 0.00100 M , phpI4z4GK.png Hg 22+ = 0.867 , phpZgVl1z.png Br - = 0.964

Consider dissociation of Hg2Br2 in KBr solution. Hg2Br2 (s) phptNicRQ.png Hg 22+ (aq) + 2  Br - (aq)

For above reaction, Ksp =( [Hg 22+ ] phpmmJYZm.png Hg 22+ ) ( [Br - ] 2phpVGR0PQ.png2 Br - )

phpR3rnEg.png 5.6 php0GlpCp.png 10 - 23 = [Hg 22+ ] 0.867 (0.00100) 2 ( 0.964) 2

5.6 phppid3jH.png 10 - 23 = [Hg 22+ ] 8.057 \times 10 -07

[Hg 22+ ] = 5.6 \times 10 - 23 / 8.057 \times 10 -07 = 7.0 \times 10 -17 M

ANSWER : [Hg 22+ ] = 7.0 \times 10 -17 M